OpenStudy (anonymous):

Solve for w in P = 2w + 2l, if P = 38 and l = 12.

OpenStudy (anonymous):

normal variables?

OpenStudy (diyadiya):

38 = 2w + 2*12 38 = 2w + 24 Can you do now?

OpenStudy (diyadiya):

I jusr substituted the values of P & l in the equation

OpenStudy (diyadiya):

*just

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok

OpenStudy (diyadiya):

did you get the answer?

OpenStudy (anonymous):

almost

OpenStudy (diyadiya):

what did you get?

OpenStudy (anonymous):

is it 100

OpenStudy (diyadiya):

How?

OpenStudy (anonymous):

Remember, just move the ones without letters to the one side and then operate on them

OpenStudy (anonymous):

38 times 2 + 24

OpenStudy (diyadiya):

38 = 2w + 24 Did you understand how i got this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2*24

OpenStudy (diyadiya):

Now we have to find w , so we subtract 24 from both sides

OpenStudy (anonymous):

then move it over

OpenStudy (anonymous):

14 and 0

OpenStudy (anonymous):

so we is 14

OpenStudy (diyadiya):

we have to isolate w 38 = 2w + 24 subtract 24 from both sides 38 -24 = 2w + 24 -24 14 = 2w Now we divide both sides by 2 14/2 =2w/2 14/2=w so w is?

OpenStudy (anonymous):

7

OpenStudy (diyadiya):

Right :)